A proton beam through a region of space where the magnetic field is a uniform value of 5.0 T directed along the positive z-axis. The protons have a velocity of magnitude 4.0 x 10⁵ m/s in the xz plane at an angle of 30° to the positive z-axis. Find the force on the portion

1
Expert's answer
2022-06-06T08:53:07-0400
"F=qvB\sin\theta\\
=1.6*10^{-19}*4.0*10^5*5.0*\sin30^\circ\\
=1.6*10^{-13}\:\rm N"